During 2017, Maria's Tennis Shop, Inc. had a cash flow to creditors of -$15,000 and the cash flow to stockholders for the year was $65,000. Suppose you also know that the firm's net capital spending for 2017 was $1,410,000 and that the firm reduced its net working capital investment by $77,000.

what was the firm's 2017 operating cash flow, or OCF?
Answer #1

Cash Flow from assets = Cash flow to creditors + Cash flow to Stock Holders = -15000 + 65000 = 50000

Cash flow from Assets = OCF + Decrease in Net Working Capital - Net Capital Spending
50000 = OCF +77,000 - 1,410,000
OCF = 50,000 + 77,000 - 1,410,000 = -1,283,000



Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well